OpenStudy (anonymous):

5. m + 8 = 22 (1 point) −30 14 −14 30 6. w – 8 = −12 (1 point) −20 −4 4 20 7. 3q = 12 (1 point) 4 9 15 36 8. q over 6=24 (1 point) 4 8 30 144 9. 3b – 9 = 15 (1 point) 8 6 2 –2 10. y over 5 + 10 = –15 (1 point) –125 –25 –5 –1 11. −3(4 − a) = 6 (1 point) −6 −2 2 6 12. 5(x + 6) = 45 (1 point) 15 3 –3 –15

OpenStudy (anonymous):

5.)14 6.)-4 7.)4 8.)144 9.)8 10.)-125 11.)2 12.)3

OpenStudy (anonymous):

5.) subtract 8 from both sides m=14 6.) add 8 to both sides w=-4 7.) divide 3 from both sides q=4 8.) multiply 6 on both sides q=144 9.) add 9 3b=24 then divide 3 b=8 10.) subtract 10 y/5=-25 multiply 5 y=-125 11.)-12+3a=6 3a=18 a=6 12.) 5x+30=45 5x=15 x=3

OpenStudy (anonymous):

thanks all you guys